A 8.0 MeV (kinetic energy) proton enters a 0.40 T field, in a plane perpendicular to the field. What is the radius of its path?
Answer #1

velocity of the proton is
v = sqrt(2*8*10^6*1.6*10^-19/1.67*10^-27) = 3.915*10^7 m/s

so the radius is
r = m*v / B*e = 1.67*10^-27*3.915*10^7 / (0.40*1.6*10^-19) m

r = 1.0215 m